find the 30th term for these numbers 2,8,16,26
Answer by richard1234(7193)

It can be seen that the numbers are in the form

a_1 = 2
a_2 = 2+6
a_3 = 2+6+8
a_4 = 2+6+8+10 etc.
By this logic,
a_30 = 2+6+8+10+...+62 = 2 + (68*29/2) = 988